# Broad-Line AGN at : The Black Hole Mass Function and a Connection with Little Red Dots

### v1

@@ 7 The BLAGN UV Luminosity Function
We compare our UV luminosity function to the Harikane et al. (2023), Maiolino et al. (2023a), and Matthee et al. (2024) JWST-detected BLAGN UV luminosity functions and find general agreement at the bright end. At moderate luminosities , our UV luminosity function lies between the higher UV luminosity functions from Harikane et al. (2023); Maiolino et al. (2023a) and the lower UV luminosity function from Matthee et al. (2024). We also plot and compare to the photometrically identified populations of LRDs from Kocevski et al. (2024) and Kokorev et al. (2024b) (grey squares); noting that while some of these are likely AGN (e.g. Greene et al., 2024; Kocevski et al., 2024), the origin of the continuum emission is still very uncertain. We also compare to the best-fit galaxy+AGN UV luminosity function at from (Finkelstein & Bagley, 2022, solid blue curve), their AGN best-fit curve where the AGN UV luminosity function was permitted to turn over (blue dashed curve) and their AGN best-fit curve where the AGN UV luminosity function was fit as a power-law to faint magnitudes (blue dotted curve). At all of our points at (where we have reasonable confidence in the completeness correction), our BLAGN+host UV luminosity function exceeds the Finkelstein & Bagley (2022) pre-JWST AGN UV luminosity function by dex.
@@ 8 Summary
We identify a population of 50 unique BLAGN (35 of which have not previously been identified) through a systematic search for H emission in the CEERS and RUBIES spectroscopic surveys.
@@
Ten of the BLAGN are also identified as LRDs using the criteria from Kocevski et al. (2024). These objects, when examined in aggregate, show broader H emission profiles and a higher fraction of broad component H emission than non-LRDs.
@@
Most BLAGN LRDs do not have red optical slopes due solely to emission lines contamination in their broadband magnitudes. Rather, the majority of these objects are intrinsically dust reddened in the rest-frame optical.
@@
We construct the BH mass function at after computing robust observational and line detection completeness corrections. This BH mass function shows strong agreement with the NIRCam WFSS based BH mass function presented in Matthee et al. (2024), indicating that our observational completeness correction is correct, though we extend to significantly lower BH masses. Our results also suggest that spectroscopic BH mass functions can be constructed without explicit knowledge of the underlying spectroscopic observation selection functions.
@@
Our BH mass function is consistent with a variety of theoretical predictions, indicating that the observed abundance of black holes in the early universe is not discrepant with physically-motivated predictions.
@@
Our BH mass function shows very little structure, and most closely resembles a power-law. This lack of structure—when compared with literature and toy black hole evolution models—indicates that any memory of the underlying BH seed distribution has been lost by redshift . Higher redshift observation of BLAGN will be necessary to directly probe the seed population of the first BHs and AGN.
@@
We construct the BLAGN UV luminosity function at using the same completeness parameters as our BH mass function. Our UV luminosity function is comparable to existing literature JWST-detected BLAGN UV luminosity functions, and demonstrates a much higher contribution to the AGN+galaxies UV luminosity function at than expected in the pre-JWST era. Our BLAGN luminosity function indicates that BLAGN account for % of the observed UV luminosity at , % at , and % at .
This work demonstrates the power of JWST to detect an unprecedented sample of 50 BLAGN at and to measure the black hole mass function over a wide dynamic range of mass () for the first time. Future NIRSpec MSA programs in Cycle 3 and beyond will further expand this sample, resulting in more precise BH mass functions and UV luminosity functions. The BH mass functions—especially when extended to redshifts by JWST studies of H detected BLAGN—may place constraints on black hole seeding models and provide a better understanding of the evolution and growth of SMBHs in the early universe. Simultaneously, the BLAGN UV luminosity function may reveal the role of BLAGN in the epoch of reionization. Moreover, this sample and future samples of BLAGN will offer new insight into the evolution of galaxies and their SMBHs over the first two Gyr of cosmic time.
We thank the RUBIES team (including PI’s A. de Graaff and G. Brammer) for their effort designing and executing this program, and for making the data publicly available. AJT thanks the members of the University of Texas at Austin Cosmic Frontier Center for insightful discussions during the development of this work. AJT acknowledges support from the UT Austin College of Natural Sciences, and AJT and SLF acknowledge support from STScI through award JWST-ERS-1345. RSE acknowledges generous financial support from the Peter and Patricia Gruber Foundation. YK thanks the support of the German Space Agency (DLR) through the program LEGACY 50OR2303. RA acknowledges support from project PID2023-147386NB-I00 and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. KI acknowledges support from the National Natural Science Foundation of China (12073003, 12003003, 11721303, 11991052, and 11950410493) and the China Manned Space Project (CMS-CSST-2021-A04 and CMS-CSST-2021-A06). The Flatiron Institute is supported by the Simons Foundation. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa.

### v2

@@ 7 The BLAGN UV Luminosity Function
We compare these UV luminosities to those measured directly from our PRISM spectroscopy (when available), and find that the EAZY-based magnitudes are magnitudes brighter than those measured directly from the PRISM data. As we only calibrate our PRISM data using JWST NIRCam data in the F150W, F200W, F277W, F356W, F410M, and F444W filters, the PRISM data are likely poorly calibrated at rest-frame 1500 Å ( m observed frame). Nonetheless, this offset is quite minor, and verifies the fidelity of our fitted EAZY-based magnitudes for use in the UV luminosity function.
@@
We compare our UV luminosity function to the Harikane et al. (2023), Maiolino et al. (2023b), and Matthee et al. (2024) JWST-detected BLAGN UV luminosity functions and find general agreement at the bright end. At moderate luminosities , our UV luminosity function lies between the higher UV luminosity functions from Harikane et al. (2023); Maiolino et al. (2023b) and the lower UV luminosity function from Matthee et al. (2024). We also plot and compare to the photometrically identified populations of LRDs from Kocevski et al. (2024) and Kokorev et al. (2024b) (grey squares); noting that while some of these are likely AGN (e.g. Greene et al., 2024; Kocevski et al., 2024), the origin of the continuum emission is still very uncertain. We also compare to the best-fit galaxy+AGN UV luminosity function at from (Finkelstein & Bagley, 2022, solid blue curve), their AGN best-fit curve where the AGN UV luminosity function was permitted to turn over (blue dashed curve) and their AGN best-fit curve where the AGN UV luminosity function was fit as a power-law to faint magnitudes (blue dotted curve). At all of our points at (where we have reasonable confidence in the completeness correction), our BLAGN+host UV luminosity function exceeds the Finkelstein & Bagley (2022) pre-JWST AGN UV luminosity function by dex.
@@ 8 Summary
We construct the BH mass function at after computing robust observational and line detection completeness corrections. This BH mass function shows strong agreement with the NIRCam WFSS based BH mass function presented in Matthee et al. (2024), indicating that our observational completeness correction is correct, though we extend to significantly lower BH masses. Our results also suggest that spectroscopic BH mass functions can be constructed without explicit knowledge of the underlying spectroscopic observation selection functions. (§6.2, Figure 12)
@@
Our BH mass function shows very little structure, and most closely resembles a power-law. This lack of structure—when compared with literature and toy black hole evolution models—indicates that any memory of the underlying BH seed distribution has been lost by redshift . Higher redshift observation of BLAGN will be necessary to directly probe the seed population of the first BHs and AGN. (§6.3, Figure 13)
@@
We construct the BLAGN UV luminosity function at using the same completeness parameters as our BH mass function. Our UV luminosity function is comparable to existing literature JWST-detected BLAGN UV luminosity functions, and demonstrates a much higher contribution to the AGN+galaxies UV luminosity function at than expected in the pre-JWST era. Our BLAGN luminosity function indicates that BLAGN account for % of the observed UV luminosity at , % at , and % at . (§7, Figure 14)
This work demonstrates the power of JWST to detect an unprecedented sample of 62 BLAGN at and to measure the black hole mass function over a wide dynamic range of mass () for the first time. Future NIRSpec MSA programs in Cycle 3 and beyond will further expand this sample, resulting in more precise BH mass functions and UV luminosity functions. The BH mass functions—especially when extended to redshifts by JWST studies of H detected BLAGN—may place constraints on black hole seeding models and provide a better understanding of the evolution and growth of SMBHs in the early universe. Simultaneously, the BLAGN UV luminosity function may reveal the role of BLAGN in the epoch of reionization. Moreover, this sample and future samples of BLAGN will offer new insight into the evolution of galaxies and their SMBHs over the first two Gyr of cosmic time.
We thank the RUBIES team (including PI’s A. de Graaff and G. Brammer) for their effort designing and executing this program, and for making the data publicly available. We thank the anonymous referee for a constructive report that helped to improve this work. AJT thanks the members of the University of Texas at Austin Cosmic Frontier Center for insightful discussions during the development of this work. AJT acknowledges support from the UT Austin College of Natural Sciences, and AJT and SLF acknowledge support from STScI through award JWST-ERS-1345. RSE acknowledges generous financial support from the Peter and Patricia Gruber Foundation. YK thanks the support of the German Space Agency (DLR) through the program LEGACY 50OR2303. RA acknowledges support from project PID2023-147386NB-I00 and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. KI acknowledges support from the National Natural Science Foundation of China (12073003, 12003003, 11721303, 11991052, and 11950410493) and the China Manned Space Project (CMS-CSST-2021-A04 and CMS-CSST-2021-A06). The Flatiron Institute is supported by the Simons Foundation. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa.
The JWST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/0eta-ym85 (catalog doi: 10.17909/0eta-ym85).